OpenStudy (anonymous):

Find the volume of a pyramid which has a square base of 10 inches and a height of 24 inches

OpenStudy (jim766):

V = 1/3 (area of base ) (hieght)

OpenStudy (jim766):

V = 1/3 (100)(24)

OpenStudy (anonymous):

how is it 100?

OpenStudy (jim766):

base is a square with a side length of 10 A = L *W A = 10*10 A = 100

OpenStudy (anonymous):

so 1/3(100)(24)

OpenStudy (anonymous):

1/3(2400)

OpenStudy (jim766):

right, keep going

OpenStudy (jim766):

multiplying by 1/3 is same as dividing by 3

OpenStudy (anonymous):

800

OpenStudy (jim766):

great! now dont forget the units in this case it will be 800 in^3 (or inches cubed)

OpenStudy (anonymous):

Thank you so much! Do you mind helping me with a similar one?

OpenStudy (anonymous):

PLease?

OpenStudy (jim766):

i will try

OpenStudy (anonymous):

A right circular cylinder has all of its dimensions multiplied by four. How is its surface area affected?

OpenStudy (anonymous):

@Jim766

OpenStudy (jim766):

hold on a second

OpenStudy (jim766):

surface area of a cylinder 2pi(r^2) + 2pi(r)(h) lets say the dimensions are r = 1 and h = 1 2pi(1^2) + 2pi(1)(1) = 4 pi

OpenStudy (jim766):

now let mult each dimension by 4 2pi(4^2) + 2pi(4)(4) 32 pi + 32pi = 64pi it went from 4 pi to 64 pi 4^3 = 64 so the answer is the surface are would be cubed